Subject: Thinfilms morning report 04/17/04
From: Jim Haydon <jhaydon@snf.stanford.edu>
Date: Thu, 17 Apr 2003 14:20:14 -0700




Balzers:    Down for water leak in RF feed through should be back up
early next week
Gryphon:    ok
Implanter:    ok
Innotec:     ok
Metalica:    ok
STS Dep:    Still working on problems hopefully they will be resolved
soon
wbgen- CTB:    ok
wbgen-LHP:    ok
wbgen-RHP:    ok
wbMETAL:    ok
wbNITRIDE:    ok
wbNONMETAL:    ok
wbDIFF:    ok
wbSILICIDE:    ok
wbGAASCL:    ok
wbMISCRES:    ok
wbOPTRES:    ok
DI WATER:    ok
NITROGEN:    ok